The length of a rectangular garden is 10 meters more than twice its width. If the area of the garden is 1000 square meters, find its dimension.
Answer by jorel555(1290)

Let w stand for width. Then length is 2w+10. So:
w(2w+10)=1000
2w²+10w-1000=0
w²+5w-500=0
(w+25)(w-20)=0
w=20,-25
The width is 20 meters, and the length is 50 meters. ☺☺☺☺